How would I solve this
thank you
Solve. 19x – 3(5x – 0.065) = 3x + 0.615

Answer by praseenakos@yahoo.com(507)

Question:
Solve. 19x – 3(5x – 0.065) = 3x + 0.615

Answer:

19x – 3(5x – 0.065) = 3x + 0.615

Remove the parenthisis using distributive property....


19x – 3*5x – 3 *-0.065 = 3x + 0.615



==> 19x - 15x + 0.195 = 3x + 0.615

==> 4x + 0.195 = 3x + 0.615

Subtract 3x from both sides....

==> 4x + 0.195 - 3x = 3x + 0.615 - 3x


==> x + 0.195 = 0.615


Subtract 0.195 from both sides...

==> x + 0.195 - 0.195 = 0.615 - 0.195

==> x = 0.42



So the solution is, x = 0.42


To check the solution, give the value x = 0.42 in the given question...then you will get the 1.875 on both sides of the equal sign.


Hope you found the explanation useful.


Regards.


Praseena.